The heating element of a clothes dryer has a resistance of 18 . A voltage of 220 V is connected across the heating element. What is the current in the heating element?
Answer #1

R = 18 Ω
V = 220 V
Let the current be I.
We know,

V = I * R
I = V/R
I = 220/18
I = 12.22 A